OpenStudy (anonymous):

For the graphed function f(x) = −(5)x − 3 + 2, calculate the average rate of change from x = 4 to x = 6. graph of f of x equals negative 1 times 5 to the x minus 3 power, plus 2 −68 −60 68 60 @satellite73 help lol

OpenStudy (anonymous):

yeah hold on let me try to interpret what you wrote

OpenStudy (anonymous):

\[f(x)=-5^{x-3}+2\]?

OpenStudy (anonymous):

\[f(x)=-1\times 5^{x-3}+2\]

OpenStudy (anonymous):

\[\frac{f(6)-f(4)}{6-4}\] is what we need to compute

OpenStudy (anonymous):

\[f(6)=-2\times 5^3+2=-25+2=-23\]

OpenStudy (anonymous):

oh not that is wrong \[f(3)=-1\times 5^3+2=-125+2=-123\] thats better

OpenStudy (anonymous):

\[f(4)=-1\times5^1+2=-5+2=-3\]

OpenStudy (anonymous):

\[f(6)-f(3)=-123-(-3)=-120\] divide by 2 get \(-60\)